OpenStudy (anonymous):

A 0.5kg piece of metal (c = 600/kgK) at 300 degree celcius is dumped into a large pool of water at 20 degrees celcius. Assuming the change in temperature of water to be negligible, calculate the overall change in entropy for the system

OpenStudy (anonymous):

we can assume that the change in temperature of water is meanless (because it is a large pool of water), so the final temperature is 20 dgree celsius. so, \[\Delta T= (300-20)K\] Then, you can determine value of Q with equation \[Q= m*c*\Delta T\] Then \[\Delta S=Q/T\] I'm stuck in here, i don't know, we must choose T in wich? May be, my brother @quarkine and @RaphaelFilgueiras can explain it anymore.

OpenStudy (anonymous):

@Rizki_Zhou pls what formula is this ΔS=Q/T

OpenStudy (anonymous):

\[\Delta S= value-of-entropy-change\]

OpenStudy (anonymous):

you can see it in http://web.mit.edu/16.unified/www/FALL/thermodynamics/notes/node41.html

OpenStudy (anonymous):

Pls guys i need help is @Rizki_Zhou in the right direction? and what is the value of T

OpenStudy (mos1635):

\[dS=dQ/T\] \[dS=d(m*c*dT)/T\] \[dS=m*c*dT/T\] \[DeltaS=\int\limits_{?}^{?}m*c*dT/T\] \[DeltaS=m*c*\int\limits_{?}^{?}dT/T\] ΔS=m*c*ln Tf/Ti that should be entropy change for the piece of metal. for the system i am not sure........

OpenStudy (anonymous):

@mos1635 wat is T final and T initial since we assume temp change is not significant

OpenStudy (mos1635):

we assume temp change is not significant for the water pool not for the piece of metal

OpenStudy (experimentx):

i think you made mathematical error.

OpenStudy (anonymous):

T final is same as that of the cold body and T initial is that of the metal before dumping...since the temp change for cold reservoir is negligible i.e its dT=0..hence entropy change is positive and it is m*c*deltaT()/T(cold) i.e negligibly small...@experimentX :what mistake did he do??

OpenStudy (experimentx):

\( \int 1/T dt = ln |T| + C\)

OpenStudy (anonymous):

integral is from ti to tf so it came out as ln (Tf/Ti)..didnt it??

OpenStudy (experimentx):

well, mathematically it should be in that way.

OpenStudy (mos1635):

\[\int\limits_{Ti}^{Tf?}(1/T)dT=\ln T (from Ti \to Tf)\] lnTf-lnTi=ln(Tf/Ti)

OpenStudy (experimentx):

also integrating gives the total entropy ... not change in entropy.

OpenStudy (anonymous):

@all im lost here..wat the final solution to this

OpenStudy (anonymous):

i think it is true as the hot body is getting colder so its entropy change is negative...since ti is more than tf,this should work fine...also we integrate both sides so there is a Sf-Si term on left side which is same as change in entropy...

OpenStudy (experimentx):

@woleraymond you know the answer??

OpenStudy (mos1635):

also integrating gives the total entropy ... not change in entropy. on the contrary my dear Watson on the contrary Q is the heat during a change not for a state so ΔS it is

OpenStudy (experimentx):

please check heat transfer between tow reservoirs http://web.mit.edu/16.unified/www/FALL/thermodynamics/notes/node41.html

OpenStudy (mos1635):

ΔS for pool ΔS=Qfrom metal/T of pool ΔS=m*C*|ΔT| of metal/T of pool and finaly ΔS total=ΔSmetal+ΔSpool

OpenStudy (anonymous):

also deltaT()/ i wrote is actually Tf-Ti itself andit wont be negligile.. @experimentX :that link which @Rizki_Zhou shared is a nice one but the thing is the metal piece is not a reservoir...

OpenStudy (experimentx):

do we really have to do |ΔT| for metal??

OpenStudy (anonymous):

oh yeah, google had answer all our question.

OpenStudy (mos1635):

give them then to @woleraymond

OpenStudy (anonymous):

@experimentX of course not...else we will lose the info about whether entropy is increasing or decreasing...

OpenStudy (anonymous):

@experimentX i dont know d answer...@Rizki_Zhou pls what is the solution

OpenStudy (experimentx):

not really sure if this works out @mos1635 was right about earlier. http://www.wolframalpha.com/input/?i=0.5+*+600+*+ln+%28293%2F573%29+%2B+0.5+*+600+*+280%2F293

OpenStudy (anonymous):

@experimentX pls clarify this equation 0.5 * 600 * ln (293/573) + 0.5 * 600 * 280/293 how do u come about 280 i also tink @mos1635 was right too

OpenStudy (experimentx):

0.5 * 600 * ln (293/573) <--- entropy change of metal piece 0.5 * 600 * 280/293 <----- entropy of sink

OpenStudy (experimentx):

first one is same as of mos the second is delQ/T provided in the link pasted above.

OpenStudy (experimentx):

I think i got around 85.5

OpenStudy (experimentx):

looks like i didn't see ln in mos equation ... sorry about that

OpenStudy (anonymous):

@experimentX 280 is 300- 20 in degrees why others are in kelvin, why?

OpenStudy (experimentx):

http://www.assignmentexpert.com/free-questions/question-on-physics-molecular-physics-and-thermodynamics-8649.html looks like it's right the difference in kelvin scale and Celsius scale is same.

OpenStudy (anonymous):

thanks @experimentX

OpenStudy (experimentx):

thanks to everyone